# Learning Multi-Index Models with Neural Networks via Mean-Field Langevin Dynamics

### v1

@@ 1 Introduction
On the other hand, neural networks can efficiently approximate arbitrary multi-index models regardless of the generative/leap exponent $s$ [8, 35]; moreover, if the (polynomial) optimization budget is not taken into consideration, there exist computationally inefficient training algorithms that can achieve sample complexity independent of $s$ [6, 44]. Intuitively speaking, a function depending on $k=O_{d}(1)$ directions of the input data has $kd=O(d)$ parameters to be estimated, and hence the information theoretically optimal algorithm only requires $n\asymp d$ samples. However, thus far it has been relatively unclear whether standard first-order optimization algorithms for neural networks inherit this optimality.
@@ 1.2 Related Works
The training dynamics of neural networks in the mean-field regime is described by a nonlinear partial differential equation in the space of parameter distributions [17, 51, 60]. Unlike the neural tangent kernel (NTK) description [40, 24] that freezes the parameters around the random initialization, the mean-field regime allows for the parameters to travel and learn useful features, leading to improved statistical efficiency. While convergence analyses for mean-field neural networks are typically qualitative in nature, in that they do not specify the rate of convergence or finite-width discrepancy, the mean-field Langevin algorithm that we study is a noticeable exception, for which the convergence rate [38, 57, 22] as well as uniform-in-time propagation of chaos [25, 62, 64, 43] have been established.
@@
The benefit of feature learning has also been studied in a “narrow-width” setting, where parameters of the neural network align with the low-dimensional target function during gradient-based training. Examples of low-dimensional targets include single-index models [7, 12, 10, 48, 31, 46] and multi-index models [30, 1, 2, 29, 9, 27, 69]. While the information-theoretic threshold for learning such functions is $n\gtrsim d$ [50, 15, 32], the complexity of gradient-based learning is governed by properties of the link function. For instance, in the single-index setting, prior works established a sufficient sample size of $n\gtrsim d^{\Theta(s)}$ where $s$ is the information exponent for one-pass SGD on the squared/correlation loss [28, 7, 10, 31], and the generative exponent [32] when the algorithm can reuse samples or access a different loss [33, 46, 3, 42]. This presents a gap between the information-theoretically achievable sample complexity and the performance of neural networks optimized by gradient descent, which we aim to close by studying the statistical efficiency of mean-field neural networks.
@@ 2 Preliminaries: Optimization in Measure Space
$\mathrm{d}\boldsymbol{w}^{t}_{j}=-m\nabla_{\boldsymbol{w}_{j}}J(\boldsymbol{w}^{t}_{1},\ldots,\boldsymbol{w}^{t}_{m})\mathrm{d}t+\sqrt{\frac{2}{\beta}}\mathrm{d}\boldsymbol{B}^{j}_{t}\quad\text{ for }\quad 1\leq j\leq m,$ (2.3)
@@
Notice that the neural network and the regularizer in (2.2) are both invariant under permutations of the weights $(\boldsymbol{w}_{1},\ldots,\boldsymbol{w}_{m})$; thus, an equivalent integral representation of these functions can be written using Dirac measures $\delta_{\boldsymbol{w}_{j}}$ centered at $\boldsymbol{w}_{j}$ as follows
@@
and allow their domain to be $\mu\in\mathcal{P}_{2}(\mathcal{W})$. Let $\mathcal{J}:\mathcal{P}_{2}(\mathcal{W})\to\mathbb{R}$ be the population risk $\mathcal{J}_{\lambda}$ or the empirical risk $\hat{\mathcal{J}}_{\lambda}$. We can equivalently state the interacting SDE system (2.3) as (see e.g. [22, Proposition 2.4])
$\mathrm{d}\boldsymbol{w}^{t}_{j}=-\nabla_{\boldsymbol{w}}\mathcal{J}^{\prime}[\mu_{\boldsymbol{W}^{t}}](\boldsymbol{w}^{t}_{j})\mathrm{d}t+\sqrt{\frac{2}{\beta}}\mathrm{d}\boldsymbol{B}^{j}_{t}\quad\text{ for }\quad 1\leq j\leq m,$ (2.5)
@@
$\!\!\mathrm{d}\boldsymbol{w}^{t}=-\nabla_{\boldsymbol{w}}\mathcal{J}^{\prime}[\mu_{t}](\boldsymbol{w}^{t})\mathrm{d}t+\sqrt{\frac{2}{\beta}}\mathrm{d}\boldsymbol{B}_{t}\quad\text{ and }\quad\partial_{t}\mu_{t}=\boldsymbol{\nabla}\cdot(\mu_{t}\nabla\mathcal{J}^{\prime}[\mu_{t}])+\beta^{-1}\boldsymbol{\Delta}\mu_{t},$ (2.7)
@@
which is referred to as the log-Sobolev inequality (LSI). If the measure ${\mathrm{d}\nu_{\mu_{t}}}\propto\exp(-\beta\mathcal{F}^{\prime}_{\beta}[\mu_{t}])\mathrm{d}\tau$ satisfies LSI with constant $C_{\mathrm{LSI}}$ for all $t\geq 0$, $\mu_{t}$ enjoys the following exponential convergence
@@ 3 Learning Multi-index Models in the Euclidean Setting
In this section, we consider learning multi-index models in the Euclidean setting. For technical reasons, we use an approximation of ReLU denoted by $z\mapsto\phi_{\kappa,\iota}(z)$ for some $\kappa,\iota>1$, which is given by $\phi_{\kappa,\iota}(z)=\kappa^{-1}\ln(1+\exp(\kappa z))$ for $z\in(-\infty,\iota/2]$ and extended on $(\iota/2,\infty)$ such that $\phi_{\kappa,\iota}$ is $C^{2}$ smooth, $\mathopen{}\mathclose{{}\left|{\phi_{\kappa,\iota}}}\right|\leq\iota$, $\mathopen{}\mathclose{{}\left|{\phi^{\prime}_{\kappa,\iota}}}\right|\leq 1$, and $\mathopen{}\mathclose{{}\left|{\phi^{\prime\prime}_{\kappa,\iota}}}\right|\leq\kappa$. Note that $\phi_{\kappa,\iota}$ recovers ReLU as $\kappa,\iota\to\infty$. Recall that we freeze the second-layer weights as $+1$. Consequently, non-negative activations can only learn non-negative functions. To alleviate this, we choose $\mathcal{W}=\mathbb{R}^{2d+2}$, and use the notation $\boldsymbol{w}=(\boldsymbol{\omega}_{1}^{\top},\boldsymbol{\omega}_{2}^{\top})^{\top}$ with $\boldsymbol{\omega}_{1},\boldsymbol{\omega}_{2}\in\mathbb{R}^{d+1}$ to denote the first and the second half of weight coordinates, and use the activation function
@@
where $\tilde{\boldsymbol{x}}\coloneqq(\boldsymbol{x},\tilde{r}_{x})^{\top}\in\mathbb{R}^{d+1}$ for a constant $\tilde{r}_{x}$ corresponding to bias, to be specified later. The above can also be seen as a 2-layer neural network with $\phi_{\kappa,\iota}$ as activation where the second-layer weights are frozen at ±$\pm 1$.
We use the neural network and the regularizer in (2.2) with weights $\boldsymbol{W}\coloneqq(\boldsymbol{w}_{1},\ldots,\boldsymbol{w}_{m})\in\mathbb{R}^{2d+2}$, and minimize the resulting empirical risk $\hat{J}_{\lambda}(\boldsymbol{W})$ via the mean-field Langevin algorithm (MFLA), which is a simple time discretization of (2.3) with the stepsize $\eta$ and the number of iterations $l>0$,
$\boldsymbol{w}^{l+1}_{j}=\boldsymbol{w}^{l}_{j}-m\eta\nabla_{\boldsymbol{w}_{j}}\hat{J}_{\lambda}(\boldsymbol{W})+\sqrt{\frac{2\eta}{\beta}}\boldsymbol{\xi}^{l}_{j},\quad 1\leq j\leq m,$ (3.2)
@@
Even though the above assumption covers a wide range of input distributions, it is mainly motivated by the Gaussian case $\boldsymbol{x}\sim\mathcal{N}(0,\boldsymbol{\Sigma})$ where $\mathopen{}\mathclose{{}\left\|{\boldsymbol{A}\boldsymbol{x}}}\right\|$ is subGaussian with norm $\|\boldsymbol{\Sigma}^{1/2}\boldsymbol{A}\|$ for any matrix $\boldsymbol{A}$. In settings we consider, we can replace the operator norm with the Frobenius norm to obtain a weaker assumption, since $\|\boldsymbol{A}\boldsymbol{x}\|$ is roughly concentrated near its mean, scaling with $\|\boldsymbol{\Sigma}^{1/2}\boldsymbol{A}^{\top}\|_{\mathrm{F}}$. Without loss of generality, we will consider a scaling where $\mathopen{}\mathclose{{}\left\|{\boldsymbol{\Sigma}}}\right\|\lesssim 1$.
@@
The effective dimension $d_{\mathrm{eff}}$ can be significantly smaller than the ambient dimension $d$, leading to particularly favorable results in the following when $d_{\mathrm{eff}}=\operatorname{polylog}(d)$. This concept has numerous applications from learning theory to statistical estimation; see e.g. [70, 72, 37, 13]. In covariance estimation, for example, the effective dimension is typically defined as $\mathop{\mathrm{missing}}{tr}(\boldsymbol{\Sigma})/\mathopen{}\mathclose{{}\left\|{\boldsymbol{\Sigma}}}\right\|$ (e.g. [72, Example 6.4]), which is equivalent to $d_{\mathrm{eff}}$ in Definition 1 provided that $\boldsymbol{U}$ lives in the top eigenspace of $\boldsymbol{\Sigma}$. However, in general, $d_{\mathrm{eff}}$ might be larger than $\mathop{\mathrm{missing}}{tr}(\boldsymbol{\Sigma})/\mathopen{}\mathclose{{}\left\|{\boldsymbol{\Sigma}}}\right\|$, which is expected as one can imagine a supervised learning setup where the variations of $\boldsymbol{x}$ provide very little information about target directions $\boldsymbol{U}$, making the target directions more difficult to be estimated. We make the following assumption on the link function in (2.1).
@@
We emphasize that the above Lipschitz condition is only local, thus allowing polynomially growing link functions $g$. We scale the Lipschitz constant with $1/r_{x}$ to make sure $y$ has a variance of order $\Theta(1)$.
@@
Let $\boldsymbol{W}^{l}=(\boldsymbol{w}^{l}_{1},\ldots,\boldsymbol{w}^{l}_{m})$ denote the trajectory of MFLA. Then, the measure $\nu_{\mu_{\boldsymbol{W}^{l}}}\propto\exp(-\beta\hat{\mathcal{J}}_{\lambda}^{\prime}[\mu_{\boldsymbol{W}^{l}}])$ satisfies the LSI (2.9) with constant $C_{\mathrm{LSI}}$ for all $l\geq 0$. For simplicity, assume $C_{\mathrm{LSI}}\geq\beta$.
@@ 3.1 Statistical and Computational Complexity of MFLA
Remark. We make the following remarks on the computational complexity of learning multi-index models.
@@ 3.2 Utilizing the Effective Dimension
$d_{\mathrm{eff}}\asymp\begin{cases}d^{1\land(2-\alpha-\gamma)}&\quad\alpha<1,\gamma<1\\ d^{1-\alpha}&\quad\alpha<1,\gamma\geq 1\\ d^{(1-\gamma)\lor 0}&\quad\alpha\geq 1\end{cases},$
@@
Figure 1: $d_{\mathrm{eff}}$ according to Corollary 6.
The scaling of $d_{\mathrm{eff}}$ (and therefore the sample complexity) is illustrated in Figure 1. We remark that the power-law assumption in (3.9) is parallel to the source condition and capacity condition in the nonparametric regression literature [26, 20], where the capacity condition measures the decay of feature eigenvalues, and the source condition measures the alignment between the target function and feature eigenvectors.
Also, based on the above result, the width and the number of iterations in Corollary 4 both become quasipolynomial in $d$ when $\alpha,\gamma\geq 1$. This corresponds to the setting where $\boldsymbol{\Sigma}$ is approximately low-rank with most of its eigenspectrum concentrated around the first few eigenvalues, and the corresponding eigenvectors are well-aligned with the row space of $\boldsymbol{U}$.

### v2

@@ 1 Introduction
On the other hand, neural networks can efficiently approximate arbitrary multi-index models regardless of the generative/leap exponent $s$ [8, 35]; moreover, if the (polynomial) optimization budget is not taken into consideration, there exist computationally inefficient training algorithms that can achieve sample complexity independent of $s$ [6, 44]. Intuitively speaking, a function depending on $k=O_{d}(1)$ directions of the input data has $kd=O(d)$ parameters to be estimated, and hence the information theoretically optimal algorithm on isotropic data only requires $n\asymp d$ samples. However, thus far it has been relatively unclear whether standard first-order optimization algorithms for neural networks inherit this optimality.
@@ 1.2 Related Works
The training dynamics of neural networks in the mean-field regime is described by a nonlinear partial differential equation in the space of parameter distributions [17, 51, 60]. Unlike the neural tangent kernel (NTK) description [40, 24] that freezes the parameters around the random initialization, the mean-field regime allows for the parameters to travel and learn useful features, leading to improved statistical efficiency. While convergence analyses for mean-field neural networks are typically qualitative in nature, in that they do not specify the speed of convergence or finite-width discrepancy, the mean-field Langevin algorithm that we study is a noticeable exception, for which the convergence rate [38, 57, 22] as well as uniform-in-time propagation of chaos [25, 62, 64, 43] have been established.
@@
The benefit of feature learning has also been studied in a “narrow-width” setting, where parameters of the neural network align with the low-dimensional target function during gradient-based training. Examples of low-dimensional targets include single-index models [7, 12, 10, 48, 31, 46] and multi-index models [30, 1, 2, 29, 9, 27, 69]. While the information-theoretic threshold for learning such functions is $n\gtrsim d$ (for isotropic data) [50, 15, 32], the complexity of gradient-based learning is governed by properties of the link function. For instance, in the single-index setting, prior works established a sufficient sample size of $n\gtrsim d^{\Theta(s)}$ where $s$ is the information exponent for one-pass SGD on the squared/correlation loss [28, 7, 10, 31], and the generative exponent [32] when the algorithm can reuse samples or access a different loss [33, 46, 3, 42]. This presents a gap between the information-theoretically achievable sample complexity and the performance of neural networks optimized by gradient descent, which we aim to close by studying the statistical efficiency of mean-field neural networks.
@@ 2 Preliminaries: Optimization in Measure Space
$\mathrm{d}\boldsymbol{w}^{t}_{j}=-m\nabla_{\boldsymbol{w}_{j}}J(\boldsymbol{w}^{t}_{1},\ldots,\boldsymbol{w}^{t}_{m})\mathrm{d}t+\sqrt{2\beta^{-1}}\mathrm{d}\boldsymbol{B}^{j}_{t}\quad\text{ for }\quad 1\leq j\leq m,$ (2.3)
@@
Notice that the neural network and the regularizer in (2.2) are both invariant under permutations of the weights $(\boldsymbol{w}_{1},\ldots,\boldsymbol{w}_{m})$; thus, an equivalent integral representation of these functions can be written using Dirac measures $\delta_{\boldsymbol{w}_{j}}$ centered at $\boldsymbol{w}_{j}$, namely
@@
with domain $\mu\in\mathcal{P}_{2}(\mathcal{W})$. Let $\mathcal{J}:\mathcal{P}_{2}(\mathcal{W})\to\mathbb{R}$ be the population risk $\mathcal{J}_{\lambda}$ or the empirical risk $\hat{\mathcal{J}}_{\lambda}$. We can equivalently state the interacting SDE system (2.3) as (see e.g. [22, Proposition 2.4])
$\mathrm{d}\boldsymbol{w}^{t}_{j}=-\nabla_{\boldsymbol{w}}\mathcal{J}^{\prime}[\mu_{\boldsymbol{W}^{t}}](\boldsymbol{w}^{t}_{j})\mathrm{d}t+\sqrt{2\beta^{-1}}\mathrm{d}\boldsymbol{B}^{j}_{t}\quad\text{ for }\quad 1\leq j\leq m,$ (2.5)
@@
$\!\!\mathrm{d}\boldsymbol{w}^{t}=-\nabla_{\boldsymbol{w}}\mathcal{J}^{\prime}[\mu_{t}](\boldsymbol{w}^{t})\mathrm{d}t+\sqrt{2\beta^{-1}}\mathrm{d}\boldsymbol{B}_{t}\quad\text{ and }\quad\partial_{t}\mu_{t}=\boldsymbol{\nabla}\cdot(\mu_{t}\nabla\mathcal{J}^{\prime}[\mu_{t}])+\beta^{-1}\boldsymbol{\Delta}\mu_{t},$ (2.7)
@@
which is referred to as the log-Sobolev inequality (LSI). If the measure ${\mathrm{d}\nu_{\mu_{t}}}\propto\exp(-\beta\mathcal{J}^{\prime}[\mu_{t}])\mathrm{d}\tau$ satisfies LSI with constant $C_{\mathrm{LSI}}$ for all $t\geq 0$, $\mu_{t}$ enjoys the following exponential convergence
@@ 3 Learning Multi-index Models in the Euclidean Setting
In this section, we consider learning multi-index models in the Euclidean setting. For technical reasons, we use an approximation of ReLU denoted by $z\mapsto\phi_{\kappa,\iota}(z)$ for some $\kappa,\iota>1$, which is given by $\phi_{\kappa,\iota}(z)=\kappa^{-1}\ln(1+\exp(\kappa z))$ for $z\in(-\infty,\iota/2]$ and extended on $(\iota/2,\infty)$ such that $\phi_{\kappa,\iota}$ is $C^{2}$ smooth, $\mathopen{}\mathclose{{}\left|{\phi_{\kappa,\iota}}}\right|\leq\iota$, $\mathopen{}\mathclose{{}\left|{\phi^{\prime}_{\kappa,\iota}}}\right|\leq 1$, and $\mathopen{}\mathclose{{}\left|{\phi^{\prime\prime}_{\kappa,\iota}}}\right|\leq\kappa$. Note that $\phi_{\kappa,\iota}$ recovers ReLU as $\kappa,\iota\to\infty$. Recall that we freeze the second-layer weights at $+1$. Consequently, non-negative activations can only learn non-negative functions. To alleviate this, we choose $\mathcal{W}=\mathbb{R}^{2d+2}$, and use the notation $\boldsymbol{w}=(\boldsymbol{\omega}_{1}^{\top},\boldsymbol{\omega}_{2}^{\top})^{\top}$ with $\boldsymbol{\omega}_{1},\boldsymbol{\omega}_{2}\in\mathbb{R}^{d+1}$ to denote the first and the second half of weight coordinates, and use the activation function
@@
where $\tilde{\boldsymbol{x}}\coloneqq(\boldsymbol{x},\tilde{r}_{x})^{\top}\in\mathbb{R}^{d+1}$ for a constant $\tilde{r}_{x}$ corresponding to bias, to be specified later. The above can also be seen as a 2-layer neural network with activation $\phi_{\kappa,\iota}$ and second-layer weights frozen at ±$\pm 1$.
We use the neural network and the regularizer in (2.2) with weights $\boldsymbol{W}\coloneqq(\boldsymbol{w}_{1}^{\top},\ldots,\boldsymbol{w}_{m}^{\top})^{\top}\in\mathbb{R}^{(2d+2)\times m}$, and minimize the resulting empirical risk $\hat{J}_{\lambda}(\boldsymbol{W})$ via the mean-field Langevin algorithm (MFLA), which is a simple time discretization of (2.3) with the stepsize $\eta$ and the number of iterations $l>0$,
$\boldsymbol{w}^{l+1}_{j}=\boldsymbol{w}^{l}_{j}-m\eta\nabla_{\boldsymbol{w}_{j}}\hat{J}_{\lambda}(\boldsymbol{W})+\sqrt{2\eta\beta^{-1}}\boldsymbol{\xi}^{l}_{j},\quad 1\leq j\leq m,$ (3.2)
@@
One example for the above assumption is the Gaussian case $\boldsymbol{x}\sim\mathcal{N}(0,\boldsymbol{\Sigma})$ where $\mathopen{}\mathclose{{}\left\|{\boldsymbol{A}\boldsymbol{x}}}\right\|$ is subGaussian with norm $\|\boldsymbol{\Sigma}^{1/2}\boldsymbol{A}\|$ for any matrix $\boldsymbol{A}$. In settings we consider, we can replace the operator norm with the Frobenius norm to obtain a weaker assumption, since $\|\boldsymbol{A}\boldsymbol{x}\|$ is roughly concentrated near its mean, scaling with $\|\boldsymbol{\Sigma}^{1/2}\boldsymbol{A}^{\top}\|_{\mathrm{F}}$. Note that Assumption 1 can cover much broader settings than Gaussianity, e.g. it is satisfied when $\boldsymbol{x}=\boldsymbol{\Sigma}^{1/2}\boldsymbol{z}$ and $\boldsymbol{z}$ has mean-zero i.i.d. subGaussian coordinates [70, Theorem 6.3.2.]. Without loss of generality, we will consider a scaling where $\mathopen{}\mathclose{{}\left\|{\boldsymbol{\Sigma}}}\right\|\lesssim 1$.
@@
The effective dimension $d_{\mathrm{eff}}$ can be significantly smaller than the ambient dimension $d$, leading to particularly favorable results in the following when $d_{\mathrm{eff}}=\operatorname{polylog}(d)$. This concept has numerous applications from learning theory to statistical estimation; see e.g. [70, 72, 37, 13]. In covariance estimation, for example, the effective dimension is typically defined as $\mathop{\mathrm{missing}}{tr}(\boldsymbol{\Sigma})/\mathopen{}\mathclose{{}\left\|{\boldsymbol{\Sigma}}}\right\|$ (e.g. [72, Example 6.4]), which is equivalent to $d_{\mathrm{eff}}$ in Definition 1 provided that $\boldsymbol{U}$ lives in the top eigenspace of $\boldsymbol{\Sigma}$. However, in general, $d_{\mathrm{eff}}$ might be larger than $\mathop{\mathrm{missing}}{tr}(\boldsymbol{\Sigma})/\mathopen{}\mathclose{{}\left\|{\boldsymbol{\Sigma}}}\right\|$, which is expected as one can imagine a supervised learning setup where the variations of $\boldsymbol{x}$ provide very little information about target directions $\boldsymbol{U}$, making estimation more difficult. We make the following assumption on the link function in (2.1).
@@
Note that the above Lipschitz condition is only local, thus allowing polynomially growing link functions $g$. We scale the Lipschitz constant with $1/r_{x}$ to make sure $y$ has a variance of order $\Theta(1)$.
@@
Let $\boldsymbol{W}^{l}=(\boldsymbol{w}^{l}_{1},\ldots,\boldsymbol{w}^{l}_{m})$ denote the trajectory of MFLA. We assume the measure $\nu_{\mu_{\boldsymbol{W}^{l}}}\propto\exp(-\beta\hat{\mathcal{J}}_{\lambda}^{\prime}[\mu_{\boldsymbol{W}^{l}}])$ satisfies the LSI (2.9) with constant $C_{\mathrm{LSI}}$ for all $l\geq 0$, and for simplicity, $C_{\mathrm{LSI}}\geq\beta$.
@@ 3.1 Statistical and Computational Complexity of MFLA
Remark. We make the following remarks on the complexity of learning multi-index models.
@@ 3.2 Utilizing the Effective Dimension
Figure 1: (a) $d_{\mathrm{eff}}$ according to Corollary 6. (b) Test loss from MFLA, details in Appendix E.
@@
$d_{\mathrm{eff}}\asymp\begin{cases}d^{1\land(2-\alpha-\gamma)}&\quad\alpha<1,\gamma<1\\ d^{1-\alpha}&\quad\alpha<1,\gamma\geq 1\\ d^{(1-\gamma)\lor 0}&\quad\alpha\geq 1\end{cases},$ (3.10)
@@
The scaling of $d_{\mathrm{eff}}$ (hence the sample complexity) is illustrated in Figure 1(a). We remark that the power-law assumption (3.9) is parallel to the source condition and capacity condition in the nonparametric regression literature [26, 20], where the capacity condition measures the decay of feature eigenvalues, and the source condition measures the alignment between the target and feature eigenvectors.
Also, based on (3.10), the width and the number of iterations in Corollary 4 both become quasipolynomial in $d$ when $\alpha,\gamma\geq 1$. This corresponds to the setting where $\boldsymbol{\Sigma}$ is approximately low-rank with most of its eigenspectrum concentrated in the first few eigenvalues, and the corresponding eigenvectors are aligned with the row space of $\boldsymbol{U}$.